OpenStudy (anonymous):

given a mean of 7.6 and a standard deviation of 3.1, what is the z-score value 5 rounded to the nearest tenth? please help ;( the answer choices are 0.8 , 0.3 , -0.3 , -0.8

jimthompson5910 (jim_thompson5910):

The z-score formula is \(\Large z = \frac{x - \mu}{\sigma}\) \(\Large \mu\) is the mean \(\Large \sigma\) is the standard deviation \[\Large z = \frac{x - \mu}{\sigma}\] \[\Large z = \frac{x - 7.6}{3.1}\] \[\Large z = \frac{5 - 7.6}{3.1}\] \[\Large z = ???\]

OpenStudy (some.random.cool.kid):

work backwards plug in the numbers in your eqaution.

OpenStudy (anonymous):

can you please give me the answer D;

jimthompson5910 (jim_thompson5910):

you just need to evaluate \(\Large \frac{5 - 7.6}{3.1}\)